If c(a-b)=a(b-c) then log(a+c)+log(a-2b+c) divided by log(a-c)equals to what?

Grade:12

1 Answers

Pratik Tibrewal
10 years ago
2ac = b(a+c)
log(a+c) + log(a-2b+c) = log((a+c)^2 - 2b(a+c)) = log((a+c)^2-4ac) = log((a-c)^2) = 2log(a-c)

when divided by log(a-c), it gives 2

Thanks and Regards,
Pratik Tibrewal,
askiitian faculty,
BTech IITG